### v1

# Collective Intelligence Outperforms Individual Talent: A Case Study in League of Legends

Abstract
11footnotetext: Corresponding author: Sajan Maharjan, School of Computing, University of Leeds, Leeds, UK, E-mail: scsmah@leeds.ac.uk
Gaming environments are popular testbeds for studying human interactions and behaviors in complex artificial intelligence systems. Particularly, in multiplayer online battle arena (MOBA) games, individuals collaborate in virtual environments of high realism that involves real-time strategic decision-making and trade-offs on resource management, information collection and sharing, team synergy and collective dynamics. This paper explores whether collective intelligence, emerging from cooperative behaviours exhibited by a group of individuals, who are not necessarily skillful but effectively engage in collaborative problem-solving tasks, exceeds individual intelligence observed within skillful individuals. This is shown via a case study in League of Legends, using machine learning algorithms and statistical methods applied to large-scale data collected for the same purpose. By modelling systematically game-specific metrics but also new game-agnostic topological and graph spectra measures of cooperative interactions, we demonstrate compelling insights about the superior performance of collective intelligence.
1 Introduction
The recent advancements in computing technology have evolved the way individuals socialize and interact; online gaming platforms being one of the most prominent mediums for social interaction and cooperation. Zhong reports that in China there are 147 million people playing online games, thus representing new forms of community, social interaction, and collaboration [1]. These online gaming environments are popular testbeds for studying human interactions and behaviours in complex artificial intelligence systems. The simulation of diverse scenarios ranging from strategic decision-making to adaptive behaviours within online gaming environments allows the testing of complex real-world mechanisms under controlled low-cost conditions. Furthermore, the virtual worlds of such online games are populated with autonomous artificial agents and other human players, thus offering opportunities for interesting interactions that guide players and enhance social dynamics [2], resulting in a collective intelligence derived from the synergistic cooperation of humans and computing agents. Collective intelligence is the general ability of a particular group to perform well across a wide range of different tasks [3]. In the field of computer science, it involves the study of how groups of one person and one computer interact (human-computer interaction) as well as how larger groups of people interact with computing technologies to cooperate and make intelligent decisions, for instance, in the development of collaborative software such as Wikipedia [4].
Collective intelligence is prominently demonstrated in many online gaming environments among which multiplayer online battle arena (MOBA) games serve as a notable example. MOBA games are a subgenre of real-time strategy games in which two teams, each consisting of 5 individual players, compete against each other with each player controlling a single character (called a champion), while strategy revolves around individual character development and cooperative team play in combat [5]. Riot Games’ League of Legends is the most popular MOBA game played worldwide with over 150 million monthly active players [6]. League of Legends features 169 uniquely playable champions with their own set of skills, that relies on an individual player’s micro-management ability, paired with a strong understanding of champion mechanics along with the collective effort of a team to think strategically and showcase team-work abilities. The game is structured around a map called Summoner’s Rift, and the primary objective of the game is to destroy the opponent team’s base, a tower structure called the Nexus, while the game progresses with champion development via earning gold, gaining experience, crafting items and scoring kills.
The collaborative dynamics observed among randomly assigned players within a team, each controlling a distinct champion with unique sets of abilities and working towards the shared objective of winning the game, potentially exemplifies the presence of collective intelligence in MOBA games. In fact, exchange of in-games chat messages, signals and assists to secure game objectives elicits a quantitative measure of collective intelligence. Leavitt et al. show that non-verbal communication in the form of in-game pings improves team strategy and performance [7]. On the other hand, individual players can excel in executing their champion mechanics and abilities, thus displaying high levels of individual intelligence. Using in-game data to model various metrics of individual and collective intelligence, this paper aims to outline the importance of such individual as well as team-based collective intelligence metrics in influencing the outcome of a game. That is, it tests the hypothesis H1: collective intelligence outperforms individual intelligence of the players in winning a game. Additionally, Riot Games maintains a two-tiered in-game ranking system (called tier and division) that establishes a hierarchy of players based on their recent in-game performances. Using such information, this paper also tests another hypothesis, that is, H2: Highly-skilled players exhibit higher levels of collective intelligence compared to less-skilled players. The main contributions of this work are – (i) a framework to evaluate individual and collective intelligence of players in League of Legends using in-game metrics and machine learning algorithms, and (ii) the availability of dataset and source code that can be used to reproduce the results.
2 Related Work
Collective intelligence can be observed in human-computer interaction, collective-decision making and collaborative software development; in the decentralized collective learning of self-managed economies [8], swarm intelligence in the large-scale coordination of drones [9], etc. Woolley et al. first provided the evidence for a collective intelligence factor, c in human groups, suggesting that c is not correlated with the average or maximum individual intelligence of the group members, but rather it emerges from the way group members interact [10]. Krause et al. conclude that the potential benefits of human swarm intelligence depend on the type of problem, while advocating for diversity in groups, and that high-performing individuals can be outcompeted by a similar size group of low-performing individuals [11]. Likewise, Santos et al. argue that social diversity catalyzes cooperative behaviour [12]. The sports and gaming industries have integrated collective intelligence paradigms to solve problems and generate insights beyond the individual capabilities of a single player.
Duch et al. quantified the contribution of individual players towards the overall team performance during the 2008 European Cup football tournament via network graphs that modeled the flow of passes among players in a team [13]. Grund confirmed that centrality of interaction between players leads to a decreased team performance by analyzing soccer passes of teams in the English Premier League [14]. Szabo et al. investigated team behavioral dynamics in escape room settings and reported that more effective teams tend to solve problems via coordinated, balanced communication patterns, maintaining a dynamic alternation between different types of tasks [15]. In the realm of digital gaming, 42 Entertainment launched I Love Bees, an alternate reality game used for the marketing of Halo 2 in which over 600,000 players collaborated in solving puzzles related to the game prior to its release [16]. Gonzalez-Pardo et al. conducted an empirical study describing how collective intelligence algorithms like ant-colony optimization and genetic algorithms can be used for problem-solving within video games [17]. Kokkinakis et al. explore the relationship between fluid intelligence and video game performance in MOBAs under controlled laboratory conditions, and found that fluid intelligence correlates significantly with in-game rank [18].
Specifically within League of Legends, Kim et al. have shown that teams exhibit group collective intelligence and that such measures for collective intelligence correlate with their in-game performance [19]. In addition, Kim et al. have shown that collective intelligence can predict the competitive performance of teams controlling for the amount of time that different individuals play together as a team, and that collective intelligence is positively correlated with the presence of a female team member and the average social perceptiveness [20]. Sapienza et al. study the evolution of individual performance within ad-hoc teams and claim that player performances in successive matches deteriorates over the course of a gaming session [21]. Kou and Gui report the existence of rich social interactions within temporary teams via semi-structured interviews with experienced players [22]. Mora-Cantallops and Sicilia have identified 4 different clusters of players associated with team cohesiveness, based on a systematic classification of player-centric networks measured by a set of graph metrics [23].
Subsequent research by Mora-Cantallops and Sicilia investigate team efficiency in professional player settings through directed graph networks that indicate the flow of assists between players and argue that team efficiency is positively affected by interactions between players while centralization of resources is detrimental [24]. Likewise, Kho et al. employ neural networks to reverse engineer games played in three professional league settings and have established an induced logic across different leagues that determine the winning outcome of a game depending on different objectives scored within the game [25]. Maymin generated high-frequency data on millions of ranked games and calibrated a live, in-game win probability model based on the conditions of the game at each moment [26]. Also, Ani et al. use ensemble methods to predict the outcome of a match, from a dataset of 1500 games using 97 different pre-match and within-game features.
The novel contribution of this research emerges from an inter-disciplinary approach that explores collective intelligence among random groups of players in the virtual world of League of Legends, by using different machine learning methods to predict the winning outcome of games. That is, the novelty of this work does not lie in the sophisticated design of machine learning models to better predict the winning outcome of a game, but rather quantifies and contrasts the importances of individual and collective intelligence behaviours among players via an empirical analysis of the collected data. The data-driven analysis presented in this work reveals an interesting insight that explains the nature of individual and collective team behaviors i.e. acquiring vs sharing and cooperative vs non-cooperative; and necessitates the importance of altruistic behaviors.
3 Methods
To evaluate the aforementioned hypotheses within the League of Legends gameplay environment, different statistical methods and machine learning algorithms are implemented. Figure 1 depicts the overall framework of the methods used in this study. Individual in-game performance metrics are extracted from the experimental dataset while team-based collective intelligence metrics are computed from their corresponding directed graph networks that outline the flow of resources between players in a team. Exploratory factor analysis is applied to both individual and team-level data, revealing latent performance factors. Such latent factors are input as features to predict the winning outcome of a game using different classification algorithms in order to understand the nature of such intelligent behaviours. Comparison of the classification results based on the individual and collective performance factors determine whether or not to accept the hypothesis H1. Furthermore, k-means clustering is used to profile players and teams into different clusters. Crosstabs analysis of the winning rate of different teams having different player clusters reinforces whether or not to accept the hypothesis H1. Finally, statistical hypothesis testing on H2 is employed to check for the significance in the difference of collective intelligence levels of teams (and players) across different ranks. Details for each step are further outlined as follows–
Figure 1: Overall architectural framework for the evaluation of collective intelligence hypotheses within League of Legends gameplay environments.
3.1 Experimental Dataset
League of Legends game data played on official servers is publicly accessible via a set of RESTful APIs provided by Riot Games [27]. In order to fetch game data for competitive matches played by a given player, a summonerName which uniquely identifies the player, is required. A preliminary list consisting of 10 random summonerNames corresponding to the 10 different ranking tiers, is used as initial seed data. For each of these seed summonerNames, the API offered by Riot Games is used to retrieve – (a) a list of $N=100$ recent matches played by the player, and for each of these matches, (b) detailed match data that include game outcome and individual match performance of the players, (c) a list of game events that occurred at periodic time intervals within the match and, (d) public rankings of the associated players. The API fetches not just the match performance of the seeded summonerName player, but also retrieves details for the other 9 players relevant in a given match. An additional set of data i.e., the most recent $N=10$ games for all the relevant players identified from the initial fetch, is collected, thus removing any initial selection bias. Using the process described above, a total of 103K matches are fetched. However, due to different modes of gameplay offered in the game, the collected dataset does not exclusively consist of ranked competitive matches only. Furthermore, two different players in this dataset may have played in the same game, or there can be games where a player was disconnected (due to Internet connection issues, or abandoning the game) from the server resulting in less than 10 players, or the game was terminated early via a remake or a surrender. To remedy such issues, the dataset was filtered to remove duplicate entries, retrieving only those games in which all 10 players were present throughout the game in a competitive ranked match mode that was not terminated early (i.e., game duration is not less than 23 minutes). Thus the experimental dataset consisted a total of 31,055 ranked matches with 164,571 unique players (see Appendix for details on data-fetch process). For each game, in-game performance metrics of all players were extracted and directed graph networks were generated to formulate team-based graph metrics.
3.2 In-game Performance Metrics
The dataset collected from the API endpoints, particularly the match detail endpoint, consists of fields that reflect the respective in-game performance of all players in a match in terms of metrics such as – experience gained, gold earned, kills score, assists given, champion deaths, vision score, elite monsters kills, towers destroyed, minions killed, etc. along with the overall match outcome. Feature scaling was performed to resolve the discrepancies in the range of values for the in-game metrics. Additionally, per-minute metrics for gold earned, assists given, vision score, experience gained and minions killed were also computed to scale and normalize the range of features in the dataset. Additionally, champions played by individual players feature their own set of unique skillshots, spells and items used during the gameplay which can impact the outcome of a game. However, incorporating such features only adds complexity without providing high-level insights on the nature of intelligent behaviors. Thus champion skillshots, spells and item builds are not accounted as in-game performance metrics for simplicity. Details on the gameplay and in-game metrics is further explained in Appendix.
3.3 Graph Metrics
Graph metrics characterize the flow and centrality of resources between individuals in a team. Such graph metrics signify how teams and individuals interact, strategize and cooperate throughout the game to secure different in-game objectives and resources. To explore the interaction among players, we utilize the match timeline data to fetch game events during which players in a team assisted one another to secure objectives and resources such as opponent kills, enemy tower destruction, elite monster kills as well as unrecorded assists for positional map-pressure resulting in a kill. Positional map-pressure refers to the intelligent positioning of a player on the map to prevent an opponent player from escaping, forcing its kill without actively participating in it (see Appendix).
Based on the flow of these assistance between players in a team, directed graph networks with 5 nodes (signifying the 5 players) with directed, weighted edges corresponding to the flow of assistance are created for both teams in each game. For any such graph network, let $N=5$ be the total number of nodes, $A$ be the total observed assists between players in the team, and $w_{ij}$ be the weight of directed edge (i.e. the number of assists) from node $i$ to node $j$. Then, the following graph metrics are computed –
Player In-Degree. This metrics signifies the weighted sum of assists received by a player from all other players in the team i.e. the in-degree of player $i$ is: $C_{ID}(i)=\sum_{j=1}^{N}w_{ji}$
Player Out-Degree. This metric signifies the weighted sum of assists given by a player to all other players in the team i.e. the out-degree of player $i$ is: $C_{OD}(i)=\sum_{j=1}^{N}w_{ij}$
Team In-Degree Centrality. This metric signifies the extent to which assists are centralized towards any player in a given team i.e.: $C_{I}=\frac{\sum_{i=1}^{N}(C_{ID}^{max}-C_{ID}(i))}{(N-1)A}$
Team Out-Degree Centrality. This metric signifies the extent to which assists are given out by any player in a given team i.e.: $C_{O}=\frac{\sum_{i=1}^{N}(C_{OD}^{max}-C_{OD}(i))}{(N-1)A}$
Effective Graph Resistance. This metric measures the robustness of a graph network [28]. A higher value of this metric signifies the lack of cooperation among nodes in the graph network. This metric is computed by extracting eigenvalues ($\mu$) from the corresponding Laplacian matrix of the graph network for each team i.e.: $EGR=N\sum_{i=1}^{N-1}\frac{1}{\mu_{i}}$
A directed graph network signifying the flow of interactions among players in a given team is illustrated in Figure 2.
Figure 2: Directed graph network for a sample team characterizing the flow of interactions between players. Each graph has 5 nodes corresponding to player roles and weighted, directed edges signify the frequency and flow of all forms of assistance from one player to another. Average values of team performance and computed graph metrics is shown.
3.4 Individual Metrics vs Collective Metrics
An individual’s performance in a game is accounted by the in-game performance metrics of the individual player such as gold earned, champion experience, kills, etc. Correlation analysis and variance inflation factor tests revealed multi-collinearity between in-game metrics which were thus removed from the feature set. Additionally, the graph metrics i.e. player in-degree and player out-degree which signify the overall receipt and provision of all assists for a single individual player is also accounted towards individual metrics. Thus, the individual metrics include – vision score per minute, gold earned per minute, champion experience per minute, player in-degree and player out-degree. The collective metrics are composed of average values of in-game performance metrics of the players in a team along with the graph metrics that signify the overall flow of assistance within the team. Thus, the collective metrics consist of average team gold earned per minute, average team vision score per minute, average team champion experience per minute, team in-degree centrality, team out-degree centrality and effective graph resistance.
3.5 Exploratory Factor Analysis
Prior work by Mora-Cantallops and Sicilia have confirmed that gold earned is one of the most important factors determining the game outcome [24]. Gold is earned by all players at a steady rate throughout the game in addition to champions and minion kills, tower takedowns and other objectives. However, only using gold earned as a feature doesn’t identify the nature of intelligent behaviors exhibited by individuals and teams. Therefore, exploratory factor analysis is performed to uncover latent factor variables that explain the nature of intelligent interactions between individuals and teams. The optimal number of explanatory factors for individuals and teams are identified via Scree analysis. Factor loading values of the latent factors infer the nature of interactions in both individuals and teams i.e. acquiring vs sharing factor and cooperative vs non-cooperative factor.
3.6 Machine Learning Algorithms
Classification: The winning outcome of a game in most cases is determined either by the exceptional performances of an individual player or by the collective efforts of a team. The explanatory factors that suggest the nature of intelligent behaviours in individuals and teams are then input as features to predict the winning outcome of the game using different classification algorithms i.e., binary logistic regression, decision tree classifier, random forest classifier, and XGBoost classifier. Classification results are evaluated in terms of accuracy, f1-score and area under the curve metrics. To accept or reject H1, the evaluation scores by the collective performance factors are compared against individual performance factors.
Clustering: Further analysis involves profiling individual players and teams based on the explanatory factors into three distinct clusters for each. K-means clustering algorithm is used to categorize individual players and teams into their respective clusters. Based on the average values of the performance factors within each cluster, we manually add labels to these clusters. The frequency of wins (or winning rate) within such distinct clusters of players and teams are then identified. Cross tabulation analysis enables how often a team cluster with a majority of particular type of players win the game. Comparison of the winning rates of such player and teams clusters via cross-tabulation analysis provides further rigour to accept or reject the hypothesis H1.
4 Results
Three main results are illustrated in this paper –
4.1 Individual Acquisition but Collective Cooperation
Exploratory factor analysis on the individual and team collective metrics reveal latent factors which relate to acquiring vs sharing of resources in individuals and cooperative vs non-cooperative behaviors in teams. The optimal number of factors for both individual and collective team performance metrics were identified to be $N=2$ via Scree analysis. Based on the factor loading values of the latent factors, such factors could be interpreted as either acquiring or sharing of resources in individuals, and as either cooperative or non-cooperative factors in teams. Table 1 outlines the factor loading values of underlying components for both individual and teams. Upon training the match dataset represented by such latent factors at both the individual and team-level to predict the winning outcome using different classification algorithms, the feature importance values for such latent factors present an interesting dynamic between individual and collective behaviour i.e., individuals attempt to win games via acquisition of resources for a player whereas teams attempt to win games via cooperative, altruistic behaviors. Figure 3 presents the feature importance values of the latent factors at the individual and collective team-level across different sets of classification algorithms.
Table 1: Factor loading values of the latent factors for individual and collective team metrics. Individual performance metrics shaded in gray benefit another player more than themselves, while non-shaded metrics are related to acquiring of resources for a player. The factor loading values for Factor 1 shows high-positive correlation with acquiring and low-negative correlation with sharing of resources, and vice-versa for Factor 2. Thus, Factor 1 and Factor 2 are distinguished as acquiring and sharing factor respectively. Likewise, collective team metrics shaded in gray relate to the resistance in the flow (or centralization) of resources and assistance within the team. The factor loading values for Factor 1 shows low positive correlation values in centralization of resources and high negative correlation for resource acquisition (average gold and experience), while Factor 2 shows higher positive correlation with metrics that signify centralization. Thus Factor 1 and Factor 2 is distinguished as cooperative and non-cooperative factor respectively.
Individual metrics Factor 1 Factor 2 Collective metrics Factor 1 Factor 2
Experience per minute 0.7226 -0.3297 Avg. experience per minute -0.8404 -0.1311
Vision score per minute -0.2679 0.9298 Team in-degree centrality 0.1518 0.4303
Player in-degree 0.7074 -0.0432 Avg. vision score per minute -0.3414 -0.1032
Player out-degree 0.0305 0.5013 Team out-degree centrality 0.0696 0.5859
Player gold per minute 0.9749 -0.2159 Avg. gold per minute -0.9837 -0.0699
Effective graph resistance 0.5320 0.5152
Figure 3: Feature importance values for individual and team collective performance factors across different classification algorithms. Across all classification models, at the individual level, the factor which represents acquisition of resources for a player is more crucial in winning the game rather than sharing. However, at the collective team level, cooperative behaviors such as resource sharing takes precedence over non-cooperative behaviors such as centralization of resources.
4.2 Collective Intelligence Outperforms Individual Intelligence
As outlined by the previous result, individuals tend to focus on the acquisition of in-game resources but cooperative behavior remains paramount at the collective team-level. Evaluation scores of different classification algorithms to predict the winning outcome based on such individual and collective team-level data reveal that collective intelligence outperforms individual intelligence. Table 2 presents the classification results of different machine learning algorithms based on the individual and collective performance factors.
Table 2: Evaluation results of different classification algorithms to predict the winning outcome based on individual and collective performance factors.
Analysis Classification Evaluation Metrics
Level Algorithms Accuracy F1-Score AUC
Individual BLR 0.7036 0.70 0.78
Decision Tree 0.6203 0.62 0.62
Random Forest 0.7115 0.71 0.78
XGBoost 0.7112 0.71 0.78
Collective BLR 0.8734 0.87 0.95
Decision Tree 0.8260 0.83 0.82
Random Forest 0.8734 0.87 0.95
XGBoost 0.8730 0.87 0.95
Furthermore, k-means clustering is used to profile teams and individual players based on these factor scores. The elbow method identified the optimal number of clusters for both individuals and teams to be $N=3$. Such clusters for individual players and teams is further labeled based on the average distribution of performance factors in each cluster. That is, individual player clusters are labeled as acquiring, sharing and average, while team clusters are labeled as cooperative, non-cooperative and average. From the experimental dataset containing 31,055 matches, there is a total of 310,550 players (10 players in each match) and a total of 62,110 teams (2 teams in each match). Table 3 shows the distribution of winning rate across different clusters of teams and individual players. Additionally, we perform cross-tabulation analysis on the distribution of winning rate in different team clusters having a majority of different types of players in Table 4. Table 5 presents a consolidated view of Table  3 and Table 4 comparing the winning rate of cooperative teams having a majority of average or sharing players (i.e. medium to low-performing individuals) with non-cooperative teams having a majority of average or acquiring players (i.e. medium to high-performing individuals).
Table 3: Distribution of winning rate across different clusters of teams and players. Cooperative teams are more likely to win games (about 89%) compared to non-cooperative ones (20%). Also, resource acquiring players are less likely to win the game than a cooperative team (i.e. 72% vs 89%). A resource sharing player has greater likelihood of winning the game than a non-cooperative team (i.e. 52% vs 20%).
Labels Games Wins Losses Win Rate
Cooperative Team 28,517 25,535 2,982 0.8954
Non-cooperative Team 26,498 5,348 21,150 0.2018
Average Team 7,095 172 6,923 0.0242
Overall 62,110 31,055 31,055 0.5000
Acquiring Player 112,862 81,672 31,190 0.7236
Sharing Player 53,413 28,077 25,336 0.5257
Average Player 144,275 45,526 98,749 0.3156
Overall 310,550 155,275 155,275 0.5000
Table 4: Cross-tabulation analysis of winning rates in cooperative/non-cooperative teams with majority ($N\geq 3$) of the acquiring, sharing or average players within the team. The rows indicates values for cooperative and non-cooperative teams whereas the column labels acquiring, sharing and average denote matches with majority of such of players in that team. Strikingly, cooperative team with a majority of resource acquiring players is the most likely to win the game.
Teams Matches Wins Winning Rate
Acquiring Sharing Average Acquiring Sharing Average Acquiring Sharing Average
Cooperative 18,079 0 2,376 17,032 0 1,831 0.9421 - 0.7706
Non-cooperative 1,366 2 18,950 747 1 2,499 0.5469 0.500 0.1318
Table 5: Winning rates of cooperative teams having a majority of low-to-medium skilled individuals compared with non-cooperative teams having a majority of medium-to-high skilled individuals. Overall, low-performing individuals in a good team can out-compete high-performing individuals in a bad team (81% vs 21%)
Team Labels Games Wins Win Rate
Coop. (T) + Low-skilled (I) 10,438 8,503 0.8146
Non-coop. (T) + High-skilled (I) 26,496 5,347 0.2018
Moreover, we compare the winning ratio of match-ups when cooperative teams with a majority of low-to-average skilled individuals face-off against non-cooperative teams having a majority of average-to-high skilled individuals as shown in Figure 4. In such matches, cooperative teams win about 98% of the games. These analyses further signify that collective intelligence outperforms individual intelligence in winning games.
Figure 4: Frequency of wins for cooperative teams having a majority of average-to-low skilled individuals and non-cooperative teams having a majority of low-to-average skilled individuals when they face each other.
4.3 Highly-skilled players exhibit higher levels of collective intelligence
[…]
Figure 5: Higher ranking teams (Diamond and above) exhibit higher levels of collective intelligence signified by lower values of effective graph resistance, in-degree centrality and out-degree centrality, compared to lower ranking teams (Silver and below).
5 Discussion
In this paper, we argue and present evidence for the statement – collective intelligence outperforms individual talent via a case study in League of Legends. By analyzing the frequency of wins, the results show that acting on the collective interest of a group leads to a higher likelihood of winning games rather than relying on individual skills and abilities. Moreover, we demonstrate that cooperative teams with low performing individuals outperform non-cooperative teams with high performing individuals.
Previous work by Mora-Cantallops and Sicilia report that centralization of resources is detrimental to team efficiency and performance [24], and our findings align with it. However, unlike this previous work in which the graph metrics – team in-degree centrality and team out-degree centrality were computed based on kills assistance, our work accounts for all forms of assists (in securing game objectives) including a tailored measure of passive assistance in positional map-pressure. Likewise, this work is the first study in which effective graph resistance is modeled to study team networks in League of Legends. Another work by Kim et. al [19] analyze group collective intelligence in controlled participants, however this work incorporates a large-scale participation of anonymous, more representative set of players, thereby enhancing the validity and general applicability of the findings. In their work, they show that higher tiered teams exhibit higher levels of collective intelligence which is congruent with our findings where we present statistically significant differences in the levels of collective intelligence across teams of different tier rankings. Additional concerns regarding the validity of our work is clarified in Appendix.
In our results, we find that non-cooperative teams having a majority of sharing (or low-performing) individuals win 50% of the games (see Table 4). However, the observed number games for such non-cooperative teams is significantly low ($N=2$). Thus, the winning rate statistic for this particular case can be misleading. To remedy such issues, a larger dataset could be collected and analyzed. As we demonstrate a dominating winning ratio of cooperative teams with low-performing individuals over non-cooperative teams with high-performing individuals in head-to-head matches, it would be interesting to identify the distribution of such winning ratio across different tier rankings. Additionally, a compelling study could be carried out whereby random, specific individuals are profiled for matches where they display exceptional performances in non-cooperative teams as well as mediocre performances in cooperative teams, and how often such displays can be attributed to wins, and their significance in climbing up the ranking ladder. As Krause et al. [11] and Santos et al. [12] advocate diversity for increased collective performance, the diversity in the types of champions picked in game, can be studied to extend this work.
The phenomenon of collective intelligence in multiplayer online gaming environments merits investigation into diverse gaming genres, for instance in First-Person Shooter (FPS) games, Massive Multi-player Online Role Playing Games (MMORPGs). Unlike FPS games (which mostly focus on individual reflex and mechanics) and MMORPGs (in which games can span over a longer time period), the authors chose to study collective intelligence in MOBA games due to their well-defined team structures, player roles, resource and objective trade-offs that bear similarities to collective actions in real-world problems. Specifically, the authors chose to conduct this study in League of Legends because – (a) it is the most popular MOBA game played worldwide, (b) features API endpoints that allow access to large-scale structured data, and (c) personally due to the authors’ familiarity and inclination towards the game. The framework that is presented in this work can also be used in other MOBA games like DOTA 2 to quantify the levels of outperformance. However, implementation will require tweaking the source code (data structures and API calls).
On a broader perspective, our results present the contradictory dynamics observed in the behavior of an individual when optimizing for self vs when optimizing for the collective. Such dynamic behaviors can also be observed in the design of sustainable infrastructure products. For instance, individuals can be tempted to use non-sustainable products which are inexpensive (immediate lower negative payoff) compared to eco-friendly products which can be costly (higher negative payoff) to implement but over a long run offers a greater benefit to the entire population [29]. Insights presented in this study can thus stipulate independent individuals to prioritize collective altruistic behaviors over individual predisposition in complex social systems. Furthermore, collective efforts can be crucial to attain the desired goals not just within online gaming ecosystems, but also in other more fundamentally critical aspects of our online (as well as offline) existence. For instance, collective efforts can help restrict the misinformation that is rampant in online social networks. Policies and regulations set up by the decentralized and collective efforts of the population in socio-economic systems such as consumer markets, governments, etc. prevent monopoly and abuse of authority resulting in products and systems that are favorable for the whole collective.
6 Conclusion
In this paper, we explored the presence of collective intelligence among players in the MOBA game, League of Legends. Our findings conclude that collective intelligence characterized by sharing of resources, flow of information and cooperative behavior outperforms individual intelligence. Our results reiterate the paradoxical imperative for collaboration and cooperation in highly competitive environments for achieving optimal outcomes. Further work complementing this research can be done by testing for the performance of collective intelligence and individual intelligence in real-world applications beyond gaming environments.
Acknowledgements
This work is funded by a UKRI Future Leaders Fellowship (MR/W009560/1): Digitally Assisted Collective Governance of Smart City Commons–ARTIO’.
Data and Code Availability
Relevant source code and sample dataset used in the analyses outlined in the paper is available at https://github.com/TDI-Lab/Collective-Intelligence-LoL.
Appendix A Appendix
A.1 League of Legends Gameplay Basics
There are different modes of gameplay available in League of Legends, for instance – ARAM, URF, Summoner’s Rift, Team Fight Tactics, etc. Ranked matches in Summoner’s Rift are the most popular and competitive modes of gameplay within the game in which two teams (each comprising of 5 players) compete to win against the other by destroying the opponent team’s base structure called nexus. Throughout the game, different types of resources and objectives are spawned, which the two teams compete for via fighting in order to earn gold and craft items which ultimately helps in destroying the nexus. Figure 6 shows a prototype map of summoner’s rift gameplay. The map comprises of three lanes – top-lane, mid-lane and bottom-lane. Accordingly, a player is assigned the role for each of the lanes. The remaining two players are assigned the roles of jungle and support. The jungler has the responsibility of moving around the lanes and assisting teammates to gain further gold via opponent kills and/or securing in-game objectives and resources. The support player accompanies the bottom-lane player (often called the Attack-Damage-Carry) providing support and utility for the player, so that over the duration of the game, the bottom-laner can scale and later ‘carry’ the game.
Figure 6: The three lanes, 2 jungles and their subdivisions. Small red and blue circles denote tower structures for the respective teams. The darker, inner-circle segment in the top-right and bottom-left corners denote each teams’ Nexus
A.2 Champion Basics
In games, each player steers independent control on in-game avatar called a champion. League of Legends features 169 unique playable champions, each with their own unique set of abilities. Such abilities can be used to deal damage, heal, shield or cast other effects to enemy or ally champions. Each champion is also equipped with 2 set of summoner spells, which provide various effects and utilities such as healing, shielding, slowing, transporting, etc. A champion requires Health Points to stay alive. A dead champion is re-spawned after a certain interval of time. Champions require mana points to cast their abilities. Items are purchased using gold. As champions gain experience and level-up, the magnitude of their abilities and base statistics rise. Figure 7 shows an in-game champion with its set of abilities, spells and item crafting.
Figure 7: Details of an in-game champion
A.3 Resources: Gold, Experience and Vision
Experience is a metric that enables players to increase their Champion’s level, in turn, increasing the Champion’s strength. Experience is primarily gained by farming minions or jungle camps. However, it is not imperative that a player kills the minion to be awarded experience. Gold is used by players to purchase in-game items, making their champion stronger. Gold is one of the most important factors which contributes towards winning the game. Gaining gold and buying items enables a champion to carry their lead forward by helping their team achieve additional objectives. Gold is earned in various ways – passively at a steady rate of time, scoring kills on enemy champions, last-hit on minions and elite-monster kills, destroying enemy tower structures, etc. Vision is a valuable resource measured by ‘vision score’. However, unlike gold or experience, no advantage is given to the player in terms of strength or currency. Players spend their gold to purchase wards – these items are placed strategically on the map to reveal dark areas that cannot be seen otherwise. These dark areas may reveal the position of an enemy player, thus, yielding a strategic advantage to the other team. Moreover, players can destroy wards placed by the opposing team – denying them such information. Doing these actions (purchasing, placing and destroying wards) increases their vision score.
A.4 Turrets and Elite-Monsters
Turrets (also known as towers), are positioned in every lane and surround each teams’ base. Turrets will attack all enemy characters within their range, one at a time. Players gain gold by attacking and destroying the opposing teams’ turrets. Destruction of an enemy turret is an example of achieving an objective. Turrets can also be destroyed by minions, in which case players receive no monetary benefit. Elite monsters are considered common objectives as they do not belong to a team, spawning mid-way at the rift. All players in the game look to kill them and reap their benefits. A small amount of gold is rewarded for securing elite monsters objectives. However, their main benefit is improving the base statistics of players (of the team that secure the objective) such as movement speed, healing rate, abilities cool-down, etc, which may be temporary or over the duration of the entire game.
A.5 Kills and Assists
A player is awarded with a kill score if the player uses their abilities or spells such that an enemy’s health points is reduced to zero. A player scoring a kill on enemy champion is rewarded with a significant amount of gold. A player scores an assist when the player helps a team-mate score a kill on an enemy champion. Assists are also awarded for tower destruction and elite-monster kills. Assistance can be in the form of dealing damage to enemy champions or providing utility effects such as healing, shielding, etc. to ally champions that results in the kill. Players receive a small amount of gold when providing assists.
A.6 Positional map-pressure assist
A new tailored metric called the “positional map-pressure" assist, which reflects how a player within a team exhibits intelligent map positioning thereby resulting in an opponent kill, is computed for each match. Figure 8 depicts how positional map-pressure can occur in-game. Opponent kill assists, tower destruction assists or elite monsters kill assists by a given player can be directly retrieved by processing the match timeline data. However, positional map-pressure assist is not directly accounted within the game. As such, by using the match timeline data, positioning of all team players an opponent kill event were identified, and a player was awarded a positional map-pressure assist if the player’s position was within a specified radius during the kill event.
Figure 8: An instance of gameplay on the Summoner’s Rift map in which an opponent kill is scored due to positional map-pressure. Blue team player A is attempting to score an opponent kill on red team player Y, while Y attempts to flee. Another blue team player B obstructs Y from escaping, thus inciting Y to engage in a fight resulting in a kill score for blue. While B does not actively assist A in the kill process, B’s positioning induces map pressure causing Y’s death. B is awarded map-pressure assist if B is positioned within a specified radius during the kill event.
A.7 Ranking System and League Points
The ranking system used in League of Legends employs a two-tiered structured ladder designed to measure and reward player skill through competitive gameplay in ranked matches. The ranking system consists of ten tiers: Iron, Bronze, Silver, Gold, Platinum, Emerald, Diamond, Master, Grandmaster, and Challenger. Each tier below Master is subdivided into four divisions (for example, Gold IV to Gold I), with IV being the lowest and I the highest within a tier.
League Points (LP) are the core mechanism that enable progression in the ranking system. Players earn LP by winning matches and lose LP upon defeat. In each sub-division of tiers below Master, LP are ranged between 0 and 100, such that upon exceeding 100 LP for a given sub-division within a tier, the player is promoted to the next immediate division (or the next tier if current division is I) and LP is reset to 0. For promotion in tiers above Master, there is no fixed rule, but rather depends on pool size of players in the upper tiers as well as higher amount of LP is required. The amount of LP gained or lost per match is influenced by several factors, including a hidden matchmaking Rating (MMR), which reflects a player’s skill level relative to their rank. Higher MMR results in larger LP gains for victories and smaller losses for defeats
A.8 Data Retrieval Process
Data on League of Legends ranked matches played on any of the official server maintained by Riot Games is freely accessible via a set of RESTful API endpoints. Figure 9 shows the overall data-retrieval process used in the study. An initial set of 10 (one for each ranking tier) random, hand-picked summoner names are stored in a text file as seed data. A python script reads this seed configuration file and makes the following API calls –
  • summonerName endpoint: This API endpoint fetches the corresponding unique identifier, called the puuid, for a given summoner name.
  • puuid/ids endpoint: This API endpoint returns a list of match identifiers for the most recent N-games (configurable) played by a player with a given puuid.
  • matchid endpoint: This API endpoint fetches the details of a given match id for each player in both teams such as - gold earned, items crafted, experience gained, kills scored, champion roles and levels, etc. Details of each match is saved as a JSON flatfile.
  • matchid/timeline endpoint: This API endpoint fetches the details of various game event (such as kills, level-ups, items purchase, etc.) as it occurs chronologically over the duration of the game with a given match id. Timeline details of each match id is saved as a JSON flatfile.
  • entries/by-summoner/summonerId endpoint: This API endpoint retrieves the rank information data for a given summoner identified by their unique puuid.
The client script thus reads the initial seed summoner names and writes match data and match timeline files in the local directory. Such files are merged and data filters, concerning game-type, duration, disconnection, etc. are applied to remove inconsistent rows.
Figure 9: Data retrieval process via RESTful APIs provided by Riot Games
A.9 Threats to Validity
There may be concerns of initial selection bias when handpicking the seed summonerNames used in the data collection process. However, we ensure that the additional fetch retrieves data for players that are not in the initial list. In 50% of the dataset, it is ensured that a single unique player has played no more than 2 games, resulting in a total of 164K unique players. Thus, any selection bias over the initially selected 10 seed summoners is unlikely to impact the results. Additionally, the collected dataset features games played exclusively within the European servers. League of Legends is played on different servers across different geographical locations i.e., America, Europe, Asia and Oceania, with notable differences in players’ pool size and skills. However, we expect data from different servers to exhibit similar pattern of results. In-game metrics such as towers destroyed, elite monster kills, enemy champion kills and minion kills are highly correlated with gold earned. Thus, these features are removed. Variance inflation test yields values of $VIF>5$ for the feature gold earned, which signifies high correlation with other features. However, it was not removed from the feature set as it is the most contributing feature for the acquiring factor. The suitability of factor analysis at the individual and collective levels yield KMO values of $0.57$ and $0.61$ respectively, which is poor but acceptable. Our results demonstrate that the winning rate for a cooperative team with a majority of resource acquiring players is as high as 94%. However this may be misinterpreted. It does not mean that the same cooperative team with the same set of high-performing individuals will win 94% of their games. This is because the Riot Games’ matchmaking algorithm ensures that as players win games, they are positioned higher in the ranking ladder, as a result they are matched with relatively more-skilled, difficult opponents when winning, and ensuring that most players roughly win 50% of their games. Finally, since the outcome of a match is a zero-sum game (no draws are applicable for a match), a team can lose not because it is bad, but the opponent has a small advantage over the team. On the other hand, a non-cooperative team can also win exploiting a blunder from the opponents. Such extreme cases have not been handled in the dataset.